40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 26 21:01:26 2019, gautam, Update, LSC, POX and POY locking 
    Reply  Thu Jun 27 14:28:12 2019, gautam, Update, LSC, POX and POY locking 
       Reply  Tue Jul 2 17:34:54 2019, gautam, Update, LSC, Irides opened up on EY table 
          Reply  Tue Jul 9 18:06:05 2019, gautam, Update, LSC, Y-arm ASS in a workable state ditherFreqs.pdftransRenorm.png
Message ID: 14738     Entry time: Tue Jul 9 18:06:05 2019     In reply to: 14720
Author: gautam 
Type: Update 
Category: LSC 
Subject: Y-arm ASS in a workable state 

The Y-arm ASS was tuned to be in a workable state. Basically, I followed Koji's recipe.

The SNR of the dither lines in the TRY and YARM control signals were checked - Attachment #1. The dither frequencies are marked with vertical dashed lines (can't figure out how to add 4 cursors in DTT so there's two in each row for a total of 4). A couple of days ago, when I was doing some preliminary checks, I found that the oscillator at 24.91 Hz caused a broadband increase in the TRY noise between DC and ~100 Hz. But today I saw no evidence of such behaviour. So I decided against changing the frequency.

The linearity of the demodulated error signals around the quadratic maxima of the TRY level was checked. I did not, however, investigate in detail the frequency-dependent offset Koji has reported in his elog. 

After this work, the TRY level is at 0.95. This is commensurate with the MC trans level being lower by ~7% relative to July 2018. Furthermore, the ASS servo is able to return to TRY~0.95 with a time-constant of ~5 seconds in response to misalignment of the cavity optics. After I investigate the X-arm ASS, I will reset the normalization for TRX and TRY.

Update 645pm: In the spirit of general IFO recovery, I re-centered the ITM and ETM oplev spots, and also the IR beam on the IPPOS QPD to mark the new input pointing alignment (the spot is slightly lower on the AS camera than what I remember). I then tweaked the XARM transmission to maximize it, and re-set the TransMon normalization. I edited the normalization script to comment out the normalizing of the TransMon QPD gains as the QPDs are in some kind of indeterminate state now. Attachment #2 shows the current status, you can also see the normalization being reset. LSC mode disabled for overnight.

Once the XARM ASS is also checked out, I propose moving back to locking the DRMI / PRFPMI configs. 

Attachment 1: ditherFreqs.pdf  8 kB  Uploaded Tue Jul 9 19:16:32 2019  | Hide | Hide all
ditherFreqs.pdf
Attachment 2: transRenorm.png  600 kB  Uploaded Tue Jul 9 19:51:21 2019  | Hide | Hide all
transRenorm.png
ELOG V3.1.3-